OpenStudy (anonymous):

A football player punts a ball. The path of the ball can be modeled by the equation y= -0.004x^2 + x +2.5, where x is the horizontal distance, in feet, the ball travels and y is the height, in feet, of the ball. How far from the football player will the ball land? Round to the nearest tenth of a foot.

OpenStudy (anonymous):

@amistre64 could you help?(:

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

"y" is the height of the ball. When the ball lands, the height will be zero. Set y = 0 in the equation and solve for x.

OpenStudy (anonymous):

@ranga okay so i got 0= -0.004x^2 + x +2.5 I don't know how to start to solve for x.?

OpenStudy (ranga):

We need to isolate x. Add 0.004x^2 to both sides. What do you get?

OpenStudy (anonymous):

@ranga 0.004x^2= x + 2.5 ?

OpenStudy (ranga):

Oh, I didn't notice the x before. Okay. This is a proper quadratic equation. Bring the x and the 2.5 over to the left hand side. They will change sign when they switch sides.

OpenStudy (ranga):

0.004x^2= x + 2.5 0.004x^2 - x - 2.5 = 0 divide throughout by 0.004. What do you get?

OpenStudy (anonymous):

@ranga it would be x^2 - x - 625?

OpenStudy (ranga):

The coefficient of x is -1. That needs to be divided by 0.004 as well. Also, add = 0 at the end.

OpenStudy (anonymous):

okay so it would be x^2 - (-250x) - 625 = 0? @ranga

OpenStudy (ranga):

yes. use quadratic formula to solve for x. brb

OpenStudy (anonymous):

@ranga okay. :)

OpenStudy (ranga):

This one does not factor nicely. So you have to use the quadratic formula to find x.

OpenStudy (anonymous):

@ranga okay. hold on a sec

OpenStudy (anonymous):

@ranga okay so i got -250 + √250^2-4(1*-625)/2(1)

OpenStudy (anonymous):

@ranga and then solved would be x≈2.475488,−252.475488

OpenStudy (ranga):

\[\Large x = \frac{ 250\pm \sqrt{(-250) ^{2} - 4(1)(-625)} }{ 2(1)}\]

OpenStudy (ranga):

Your signs are switched because you had -250 in the formula but it should have been +250. x ≈ -2.475488 or 252.475488. Throw away x ≈ -2.475488 (extraneous solution). So x ≈ 252.475488 They asked you to round the answer to the nearest tenth of a foot. So x = 252.5 feet.

OpenStudy (anonymous):

@ranga okay thank you so much!! <3 lifesaver.

OpenStudy (ranga):

You are welcome.